40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 242 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  5164   Wed Aug 10 02:29:38 2011 JenneUpdateSUSETMY exploration

Quote:

- Jenne will make a better kick/free-swing test later.

 02:27am, ran the new freeswinging-ifo.csh script.  It's just a copy of freeswinging-all.csh, but it doesn't include the MC mirrors, since Suresh and Kiwamu are still working.  

Now we have copies of the script for -all, -mc, -ifo to cover the various sections of the suspended interferometer.

  5163   Wed Aug 10 01:40:40 2011 KojiUpdateSUSETMY exploration

[Rana Koji Jenne Jamie]

- The situation of the ETMY suspension is improved.
- The damping servos except for Pitch are now functional.
- We intentionally turned off the damping servos for the matrix measurements.


- Opened the light door of the ETMY chamber.

- We set up the CDS SUS lockin:

        Excite UL/UR/LL/LR equally [by setting the output matrix (1, 1, 1, 1, 0)] at 3.12Hz with 2000 cnts
        Put the OSEM PD outputs into lockin one by one. For the image rejection, 0.1Hz 4th order LPF has been used though we like to use a faster settling LPF.

- Found only UL was responding to the excitation. After fitzing with the cables and connectors, it was found that the DAC card was loose from the bus.
  By pushing the card the responses have been back. [Note we had the reboot of c1iscey almost at the same time.]

- Checked the response in the I channel of the lockin.
        UL -8ish / UR +7ish / LR +5ish / LL +2ish

- Tweaked LL sensor to get better response ==> in vain. Decided to move the lower OSEM plate for the better positioning of the LR/LL.
- Got reasonable (+5ish) response for LL.

- Confirmed that the POS/YAW/SIDE damping works with positive gains. PITCH did not work with the negative gain (but that could be a good sign.)

- Let the suspension freely swinging for a while (~30min). Checked the side/pos separation. They are not perfect but seemed diagonalizable.

- Closed the light door.

- Jenne will make a better kick/free-swing test later.

  5162   Wed Aug 10 00:21:10 2011 jamieUpdateCDSupdates to peakFit scripts

I updated the peakFit routines to make them a bit more user friendly:

  • modified so that any subset of optics can be processed at a time, instead of just all
  • broke out tweakable fit parameters into a separate parameters.m file
  • added a README that describes use

These changes were committed to the 40m svn.

  5161   Wed Aug 10 00:11:39 2011 jamieUpdateSUScheck of input diagnolization of ETMX after OSEM tweaking

Suresh and I tweaked the OSEM angles in ETMX yesterday.  Last night the ETMs were left free swinging, and today I ran Rana's peakFit scripts on ETMX to check the input diagnolization:

ETMX.png

It's well inverted, but the matrix elements are not great:

       pit       yaw       pos       side      butt
UL    0.3466    0.4685    1.6092    0.3107    1.0428
UR    0.2630   -1.5315    1.7894   -0.0706   -1.1859
LR   -1.7370   -1.5681    0.3908   -0.0964    0.9392
LL   -1.6534    0.4319    0.2106    0.2849   -0.8320
SD    1.0834   -2.6676   -0.9920    1.0000   -0.1101

The magnets are all pretty well centered in the OSEMS, and we worked at rotating the OSEMS such that the bounce mode was minimized.

Rana and Koji are working on ETMY now.  Maybe they'll come up with a better procedure.

  5160   Tue Aug 9 19:53:56 2011 NicoleSummarySUSWeekly Summary

This week, I have finished assembling everything I need to begin shaking. I built an intermediary mounting stage to mount the TT suspension base to the horizontal sliding platform, finished assembling the second photodiode, finished assembling the photosensor circuit box, and calibrated the two photosensors. Today I built a platform/stage to mount the photodiodes so that they are located close enough to the mirror/suspension that they can operate in the linear range.  Below is an image of the set-up.

entiresetup.jpg

The amplifer that Koji fixed is acting a bit strange again...It is sometimes shutting off (Apparently, it can only manage to do short runs ~ 1minute? That should be enough time?).

The set-up is ready to begin taking measurements.

  5159   Tue Aug 9 16:54:47 2011 steveUpdateVACscratch on vac door

Jenne found a big scratch on the north vac door of ITMY. Fortunately it does not reach the inner annulos 0 -ring seal.

This is precisely what we have to avoid to preserve our vacuum system!

 

Attachment 1: P1080152.JPG
P1080152.JPG
  5158   Tue Aug 9 16:40:12 2011 steveUpdateGeneralclean room coat counts

I measured the particles coming off of new- unused clean room coat. Tyvek, Convertors, Allegiance #9393 measured  10 counts of 0.3 micron at 1 minute and 0 count at 7 minute.

The 0.5 micron size  measured 0 at both times. Jenne's used coat measured 20 counts of 0.3 micron  and 0 counts for 0.5 micron at 1 minute. ( counts / cf- min)

The HEPA tent background is consistently 0 when it's CP STAT 100 curtains are closed.

Attachment 1: P1080155.JPG
P1080155.JPG
  5157   Tue Aug 9 16:21:59 2011 Manuel, IshwitaUpdateWienerFilteringFirst results of our simulations

We did the simulation of the stacks by defining a transfer function for one stack (green plot) and another similar transfer function for the other stack.

We simulated the ground motion by filtering a white noise with a low pass filter with a cutoff frequency at 10Hz. (blue plot) (the ground motion for the 2 stacks are completely uncorrelated)

We simulated the electronic white noise for the seismic measurements. (black plot)

We filtered the ground motion (without the measurements electronic noise) with the stack's transfer function and subtracted them to find the mirror response (red plot), which is the target signal for the wiener filter.

We computed the static wiener filter with the target signal (distance between the mirrors) and the input data (seismic measurements = ground motion + electronic noise).

We filtered the input and plotted the output (light blue plot).

We subtracted the target and the output to find the residual (magenta plot).

We didn't figure out why the residual is above the electronic noise only under ~6hz. We tried to increase and decrease the electronic noise and the residual follows the noise still only under ~6Hz.

It also shows that the residues are above the target at frequencies over 20Hz. This means that we are injecting noise here.

simu1.png

We tried to whiten the target and the input (using an high pass filter) to make the wiener filter to care even of higher frequencies.

The residues are more omogeneously following the target.

We also plotted the Wiener filter transfer function without making whitening and with making whitening. It shows that if we do whitening we inject no noise at high frequency. But we loose efficency at low frequencies.

simu2w.png

TFwhitened2.png

We shouldn't care about high frequency, because the seismometers response is not good over 50Hz. So, instead of whitening, we should simply apply a low pass filter to the filter output to do not inject noise and keep a good reduction at low frequencies.

 

  5156   Tue Aug 9 16:00:58 2011 JennyUpdatePSLAmplitude response of PZT

AMresponsePZT.png

The top plot shows a sweep from 10 kHz to 5 MHz of the ratio of the voltage output of the PD detecting power from the NPRO laser beam and the RF source voltage (the magnitude of the complex transfer function). The black trace was taken with the laser beam blocked. For runs 2 and 3 I changed the laser temperature set point by 10 mK and 100 mK respectively to see if there was a significant change in the AM response. The bottom plots shows runs 2 and 3 compared to run 1 plotted in dB (to be explicit, i'm plotting 10 times the base 10 log of the magnitude of the ratio of two complex transfer functions). Changing the temperature seems to have only a minor effect on the output except at around 450kHz, where the response has a large peak in run 1 and much smaller peaks in runs 2 and 3. 

The traces in the top plot consist of 16 averages taken with a 300Hz IF bandwidth, 15 dBm source power (attenuated with a 6 dB attenuator) and with 20dB attenuation of the input power from the PD.

Next I'm going to probe a narrow band region where the response is low (2.0MHz or 2.4MHz perhaps) and choose a bandwidth for the dither frequency for the PDH locking.

Attachment 1: AMresponsePZT.png
AMresponsePZT.png
  5155   Tue Aug 9 15:14:46 2011 KojiUpdateGeneralin-vac work plan today

This morning Kiwamu and I have aligned the MC. Kiwamu aligned the last steering (on the OMC table) to recover the touch last week.
Then I have aligned the MC with MC1 and MC3 as the last steering did not help to get TEM00.

Before

C1:SUS-MC1_PIT_COMM = 2.6587
C1:SUS-MC1_YAW_COMM = 2.7471
C1:SUS-MC2_PIT_COMM = 3.486
C1:SUS-MC2_YAW_COMM = -1.1592
C1:SUS-MC3_PIT_COMM = -1.876
C1:SUS-MC3_YAW_COMM = 1.2829

After

C1:SUS-MC1_PIT_COMM = 2.7596
C1:SUS-MC1_YAW_COMM = 2.6627

C1:SUS-MC2_PIT_COMM = 3.486
C1:SUS-MC2_YAW_COMM = -1.1592
C1:SUS-MC3_PIT_COMM = -1.697
C1:SUS-MC3_YAW_COMM = 1.2901

Quote:

We will move on to the vertex region today.

The goal of the vertex region work is to get the pick-off beams out the chambers, including POX/Y and POP.

The work will be in parallel to the ETM woks.

The first step will be : lock and align MC with the IR beam.

 

  5154   Tue Aug 9 13:34:40 2011 NicoleSummarySUSNew Calibration Plots for Photosensors

Here are the new calibration plots for my photosensors. These calibrations were done using a translation stage.

The linear region for the first photosensor appears to be between 15.2mm and 30 mm

ps1.jpg

The linear region for the second photosensor appears to be between 12.7mm and 22.9mm

ps2.jpg

The slope for both is -0.32 V/mm  (more precisely, -0.3201 V/mm for PS 1 and -0.3195 V/mm for PS 2)

 

  5153   Tue Aug 9 11:33:33 2011 kiwamuUpdateSUSRe: ETMX free swinging data

I believe that the 17 Hz broad structure on SIDE is just because of a bad rotational angle of the SIDE OSEM.

The same structure had been observed on the EMTY_UR, and the structure became narrower after we repositioned/rotated the OSEM yesterday.

My guess is that the SIDE OSEM is now in a place where the OSEM is quite sensitive to the bounce mode

and creating the broad structure due to a bi-linear coupling between the bounce mode and low frequency signals.

Quote from #5150

There is something defintely wrong with the side sensor.  It might be the electronics as it also has this problem with it slow channel readings (my previous elog today). 

  5152   Tue Aug 9 10:15:25 2011 kiwamuUpdateGeneralin-vac work plan today

We will move on to the vertex region today.

The goal of the vertex region work is to get the pick-off beams out the chambers, including POX/Y and POP.

The work will be in parallel to the ETM woks.

The first step will be : lock and align MC with the IR beam.

  5151   Tue Aug 9 03:05:05 2011 kiwamuSummaryGeneralweekly report

Summary of the week ending Aug 8th.  Number of elog entries = 56


- VAC
 + The vent started Wednesday morning
 + Repositioning of the green periscopes and associated mirrors are done.
 + Got both of the green beams coming out from  the chambers
 + Moved the ETMX suspension tower by -8.09 inch (away from vertex)
 + Fixed the alignment of the ETMX CCD mirrors
 + Recovered the X green beam axis for the latest ETMX position

- SUS
 + oplev centered prior to the vent

- LSC
 + ETMY_TRANS_QPD didn't respond at all, needs to be fixed
 + Old MZ PD (InGaAs 2mm, @29.5MHz) has been modified for REFL33.
  The 11MHz notch circuit is at the amp side instead of the diode side. This is ready for the installation
 + REFL165 PD has been made from the old 166MHz PD.

- IOO
 + IPPOS has been sick since 19th of July, 2011
 + IPANG is clipped on a pick-up mirror on the ETMY table. QPD itself is healthy.
 + The spot positions on the MC mirrors were measured prior to the vent.
   The results are almost the same as before within a few percent difference expect for the MC2 yaw.
 + An attenuator, consisting of two HWPs and a PBS, has been installed on the PSL table for the MC low power state.
 + a 10% BS in front of the MCREFL_RFPD was replaced by a perfect reflector for the low power mode.
 + The incident power for MC was decreased to 20 mW
 + The beam axis going to MC was misalgned due to the attenutor.
    Then the beam was aligned by touching two steering mirrors on the PSL table
 + MC is able to be locked in air. The reflection DC goes from 1.4 to 0.13 V when the MC is locked.

- ABSL
 + With the mass-kicking technique, the arm lengths were measured.
   Xarm =  37.5918 m, Yarm = 37.5425 m.

- Green locking
 + Y green beam is aligned to the Y arm
 + Locking of the Y green is not robust, it needs to be revisited

- OAF
 + Wiener Filtering was applied on the data collected from the X-arm for a duration of 1500 seconds.

- Misc.
 + The hazardous waste people are moving chemicals around outside our door, and have roped off our regular front door.
 + The horizontal trolley drive of the east end crane stopped working. It will be fixed.

  5150   Tue Aug 9 02:44:32 2011 SureshUpdateSUSETMX free swinging data

I switched off damping to the ETMX and used a reduced version of freeswing-all.csh script (called freeswing-ETMX.csh) to set it swinging.    After about an hour I used the saved template ETMX/2008.08.06.xml to obtain the following plot.

 ETMX_freeswing.png

 

There is something defintely wrong with the side sensor.  It might be the electronics as it also has this problem with it slow channel readings (my previous elog today).

 

 

  5149   Tue Aug 9 02:34:26 2011 JennyUpdatePSLPZT transfer function measurement

Using a PDA255 on the PSL table, I measured the amplitude response of the NPRO PZT, sweeping from 10kHz to 5 MHz.

I took a run with the laser beam blocked. I then took three runs with the beam unblocked, changing the temperature of the laser by 10 mK between the first two runs and by 100mK between the second and third runs.

At the end of the night I turned off the network analyzer and unplugged the inputs. I'm leaving it near the PSL table, because I'd like to take more measurements tomorrow, probing a narrow bandwidth where the amplitude response is low.

On the PSL table, I'm still monitoring the reflected light from the cavity and the transmitted light through the cavity on the oscilloscope. I'm no longer driving the NPRO temperature with the lock-in.

I closed the shutter on the NPRO laser at the end of the night.

I'll log more details on the data tomorrow morning.

  5148   Tue Aug 9 02:27:54 2011 Ishwita , ManuelUpdatePEMPower spectra and Coherence of Guralps and STS2

We did offline wiener filtering on 3rd August (Elog entry) using only Guralps' channels X and Y.

Here we report the Power spectrum of the 3 seismometers (Guralp1, Guralp2, STS1) during that time.

and also the coherence between the data from different channels of the 3 seismometers.

We see that the STS is less correlated with the two Guralps. We think it is due to the wrong alignment of the STS with the interferometer's axes.

We are going to align the STS and move the seismometers closer to the stacks of the X arm.

 Pw_gurs_correct.png

 

 


 Pw_gur1_sts1_correct.png


 

 


 

 Coher_gur1_gur2_BW0.01.png

 

 



Coher_gur1_sts1_BW0.01.png

 

 



Coher_sts1_gur2_BW0.01.png

  5147   Tue Aug 9 02:03:16 2011 kiwamuSummarySUSsummary of today's work on ETMY

[Rana / Jenne / Kiwamu]

The ETMY suspension tower is currently sitting on the north side of the table for some inspections.

The adjustment of the OSEMs is ongoing.

 

(What we did)

  + Taken out two oplev mirrors, Jamie's windmill and a lemo patch panel.

  + Put some pieces of metal as makers for the original place

  + Put some makers on the distance of  dLY = -25.49 cm = -10.04 inch from the original place (see the 40m wiki).

     The minus sign means it will move away from the vertex.

  + Brought the ETMY suspension tower to the north side to do some inspections

  + Did some inspections by taking the noise spectra (#5141)

  + Adjusted the OSEM range and brought the magnets on the center of the OSEM holders by rotating and translating the OSEMs

  + During the work we found the proper PIT and YAW gains were about -5, which are the opposite sign from what they used to be.

  + Trying to minimize the cross couplings

JD: There is still some funny business going on, like perhaps the LR magnet isn't quite in the OSEM beam.  We leave the optic free swinging, and will continue to investigate in the morning.

  5146   Tue Aug 9 01:35:45 2011 SureshUpdateSUSETMX Side Sensor slow channel down for a long time

The slow signal from the side sensor on ETMX was last seen in action sometime in May 2010!  And then the frame builder has no data for a while on this channel.  After that the channel shows some bistability starting Sept 2010 but has not been working.  The fast channel of this sensor  (C1:SUS-ETMX_SDSEN_OUTPUT) does work so the sensor is working.  Probably is a loose contact... needs to be fixed.

ETMX-SDSEN_trend1.png         ETMX_SDSEN_trend2.png

 

  5145   Mon Aug 8 22:12:58 2011 NicoleSummarySUSDaily Summary

Today I balanced the mirror, finished putting together the second photosensor, and finished my photosensor circuit box! 

Upon Jamie's suggestion, I have used a translation stage to obtain calibration data points (voltage outputs relative to displacement) for the new photosensor and for the first photosensor.

I will plot these tomorrow morning (too hungry now > < )

 

Here is a photo of the inside of my circuit box! It is finally done! It is now enclosed in a nice aluminum casing ^ ^

 

frontview.jpg

Attachment 1: frontview.jpg
frontview.jpg
  5144   Mon Aug 8 20:23:14 2011 JennyUpdatePSLNetwork analyzer and PD set up to measure amplitude response of PZT

Quote:

Quote:

Today I placed the PDA255 photodiode on the PSL table to catch the small amount of beam power reflected by the second polarizing beam splitter in my setup. I plugged the PD output to the oscilloscope to measure the voltage output and positioned the PD such that the voltage output was maximized. At best I was able to achieve a 300 mV DC output voltage from the PD, (which seems a bit low, as the PD is specified to go from 0 to 5 V and the specifications say that the response becomes nonlinear after 10 mW/cm^2 and my beam has an intensity of approximately 5 mw/cm^2. I would therefore expect to get more beam power but after over an hour of maneuvering, 300 mV was the highest voltage output I could get).

I am planning, tomorrow afternoon, to take a measurement of the amplitude response of the PZT driving the NPRO laser. I moved the 4395 spectrum/network analyzer to near the PSL table and connected the RF output to an RF splitter. I fed one output of that into the PZT and the other output into the R port on the network analyzer. I fed the PD output into the A port. I plan to measure A/R as a function of driving frequency, sweeping from 10 Hz to 30 mHz.

I also worked to improve the mode matching of the NPRO beam coming from the AP table to the reference cavity. I drove the temperature of the NPRO at 0.100 Hz with an amplitude of 0.300 V, which Koji told me corresponds to a 1GHz change in the laser frequency. The transmission from the cavity is being monitored by a camera connected to a TV monitor, and also by a PD connected to an oscilloscope. I then repositioned the second lens in my mode matching setup in an attempt to increase the transmission peaks from the zeroth order spacial mode and decrease the transmission peaks from higher order modes. I may have improved the mode matching slightly but I was unable to improve it significantly.

The ABSL beam had been blocked so that it wouldn't enter the interferometer. I moved the block so that the beam I've been using is unblocked by the beam going to the interferometer is still blocked.

I positioned a fast lens (f=28.7mm) a little over an inch in front of the PDA255 in order to decrease the spot size incident on the PD. I adjusted the rotation angle of the half wave plate to maximize the transmitted power through the PBS to the cavity and minimize the power reflected to my PD. I then adjusted the lens potion to fix the beam on the PD. The voltage output of the PD is now 150mW, but I have the ability to increase the incident power by rotating the wave plate slightly.

Now all I need is to set up the network analyzer again to record the amplitude response to modulating the PZT from 10 Hz to 30 MHz, reduce the input voltage into the analyzer using a DC block.

 I rolled the network analyzer over to the PSL table (on the south side). I'm borrowing the DC block from Kiwamu's green locking setup. I'm going to first measure the amplitude response of a low pass filter to made sure that the analyzer is outputting what I expect. Then I will measure the laser PZT amplitude response. I plan to finish the measurement and return the network analyzer to it's usual location tonight.

  5143   Mon Aug 8 19:45:27 2011 jamieUpdateCDSactivateDQ script run; SUS channels being acquired again

> Also the BS is missing its DAQ channels again (JAMIE !) so we can't diagnose it with the free swinging method.

I'm not sure why the BS channels were not being acquired.  I reran the activateDQ script, which seemed to fix everything.  The BS DQ channels are now there.

I also noticed that for some reason there were SUS-BS-ASC{PIT,YAW}_IN1_DQ channels, even though they had their acquire flags set to 0.  This means that they were showing up like test point channels, but not being written to frames by the frame builder.  This is pretty unusual, so I'm not sure why they were there.  I removed them.

  5142   Mon Aug 8 15:52:39 2011 steveUpdateVACback ground rga scan

The RGA region is beeing monitored during the vent. This will tell us how clean the RGA itself is.

 

Attachment 1: bg-d5.jpg
bg-d5.jpg
Attachment 2: presBg-d5.jpg
presBg-d5.jpg
  5141   Mon Aug 8 15:20:49 2011 kiwamuUpdateSUSinspection on ETMY (round 1)

Since ETMY have been showing some strange behaviors we deeply inspect the ETMY suspension.

Here is a brief review of the ETMY suspension and a brief status update of the inspection so far.

The inspection is still ongoing.

 

(Review : How wrong ?)

First of all, let us summarize what were the observed phenomena on the ETMY suspension :

   1. unknown low frequency noise covering frequency range from 0.1 to 3 Hz in all of four face sensors (#5025, #5029).

   2. LR sensor showed a very broad bounce peak at ~ 17 Hz (#5025).

   3. The sign of some of the sensors are flipped.

   4. The control gains had to be higher than those of ETMX by 10-100 (#5025).

 

(Status update : Noise spectra)

Currently the ETMY suspension is sitting on the north side of the table for the inspection.

We took dark noise of the OSEMs when the OSEMs were taken off from the tower and put on the table.

The plot below is an example of the LR sensor spectra. Note that the whitening filters have been always ON.

Screen_shot_2011-08-08_at_15.08.23.png

The black curve is the dark noise when the sensor was off from the tower.

The blue curve is the free swinging spectrum newly taken today.

The red curve is the free swinging spectrum (damped spectrum ?) on 25th of July, this was still in vacuum.

 

The dark noise is below the free swinging spectrum from 0.2 - 30 Hz, which looks reasonable

The most interesting thing is that the free swinging spectrum became better in low frequency (below 3 Hz)

from the one measured in vacuum.

It needs more investigation to answer the reason why it happened.

Note that before we moved the tower to the current position, we looked at the OSEM-magnets relations, and found nothing was touching.

 

  5140   Mon Aug 8 14:21:03 2011 steveUpdatePEMParticle counts controlled

Quote:

[Jenne, Kiwamu]

While Kiwamu was finalizing the X green alignment, I started to prepare to remove the ETMY door, and begin checking out its OSEMs, etc, so we could start moving it to it's new place, and figure out why it's been wonky for a while.  I ran the particle counter, and we have a factor of ~5 more particles than normal.  Kiwamu and I agreed not to open ETMY.  Since we had briefly opened the IOO and Output Optics chambers to check the X green's position on the PSL table, we immediately shut those doors.  They were probably open for ~15 minutes or so.  (Yes Steve, we should have checked before opening any doors, but at least we remembered to check at all, and the doors were only open for a few minutes rather than for a few hours.)

I attach a 24hrs trend of the particle counts, for reference.  It looks like it's been a little high for a while, but today it's really dirty in the air.

 The east end particle count is 155K for 0.3 micron and 15K for 0.5 micron  counts/cf min.  In order to minimize diffusion of dirt into the IFO we set up the mobile HEPA with CP Stat 100 tent.

This set up gives us practically ZERO inside the tent

Attachment 1: P1080149.JPG
P1080149.JPG
Attachment 2: P1080150.JPG
P1080150.JPG
  5139   Mon Aug 8 13:24:00 2011 steveMetaphysicsTreasure Monday morning

Quote:

Just a reminder that a film crew will be here Monday morning, filming Christian Ott for some Discovery channel show.

They are slated to be here from 8am to 12:30pm or so.  They will take a couple of shots inside the lab, and the rest of the filming should be of Christian in the control room (which they will "clean up" and fit with "sexy lighting").  I will try to be here the whole time to oversee everything.

Konecrane Fred was early this morning. He diagnosed the ETMY crane horizontal drive gear box dead and left just before the film crew showed up.

New gear box should be here by the end of this week for installation.

The lab air quality is high ~20,000 counts of particles of 0.5 micron.  Keep an eye on this before you open the chamber.

Attachment 1: P1080147.JPG
P1080147.JPG
  5138   Mon Aug 8 12:08:05 2011 ranaUpdateSUSsus at atm

Untitled.png

A plot showing that the daily variation in the OLs is sometimes almost as much as the full scale readout (-1 to +1).

  5137   Mon Aug 8 00:58:26 2011 ranaUpdateCDSdiagonalization of MC input matrix

Besides the purpose of correctly tuning the suspensions, my hidden goal in the input matrix diagonalization has been to figure out what the 'true' sensing noise of the OSEMs is so that we can accurately predict the noise impact on the OAF.

The attached plot shows the DOFs of ITMX calibrated into microns or microrad as per Jamie's ethereal input matrix calculations.

The main result is in the ratio of POS to BUTTER. It tells us that even at nighttime (when this data was taken) we should be able to get some reduction in the arms at 1 Hz.

Whether we can get anything down to 0.1 Hz depends on how the arm control signal compares to the POS signal here. I leave it to Jenne to overlay those traces using a recent Arm lock.

Attachment 1: null.png
null.png
  5136   Mon Aug 8 00:12:58 2011 ranaUpdateCDSdiagonalization of MC input matrix

I've finally completed the SUS/peakFit/ scripts which find the new input matrix for the SUS. MC1, MC2, MC3, and ITMX have been matrix'd.

I tried to do the BS, but it came out with very funny matrix elements. Also the BS is missing its DAQ channels again (JAMIE !) so we can't diagnose it with the free swinging method.

To continue, we have to get some good data and try this again. Right now there are some weird issues with a lot of the optics. I've also set the damping gains for the optics with the new matrices.

 Ex.

new_matrix = findMatrix('ITMX')

writeSUSinmat('ITMX', new_matrix)

writeSUSinmat.m

this script writes the values to the MEDM input SUS matrix. To do the writing, I used the low level 'caput' command instead of ezcawrite since the ezca libraries are getting deprecated.

caput doesn't really have good diagnostics, so I use matlab to check the return status and then display to the terminal. You can just rerun it if it gives you an error.

 

 

A coupled of normalization notes:

1) The POS/PIT/YAW rows are scaled so that the mean of abs(FACE elements) = 1. Previously, I had the max element = 1.

2) The SIDE row is scaled so that the SIDE element = +1.

3) I then normalized the ROWS according to the geometrical factors that Jamie has calculated and almost put into the elog.

 

All these scripts have been added to the SVN. I've removed the large binary data files from the directory though. You can just rsync them in to your laptop if you want to run this stuff remotely.

  5135   Sun Aug 7 18:41:12 2011 JenneUpdateGreen LockingX green beam re-aligned

Quote:

The ETMY suspension tower is in a good place.

Quote from #5127

Kiwamu will work on the green alignment over the weekend.  Assuming everything works out, we'll try the same procedure on ETMY on Monday.

 

 You meant ETMX, right?  ETMY still hasn't been touched.

KI : sorry, I meant ETMX. I fixed the entry.

  5134   Sun Aug 7 14:11:53 2011 JenneUpdatePEMParticle counts through the roof

[Jenne, Kiwamu]

While Kiwamu was finalizing the X green alignment, I started to prepare to remove the ETMY door, and begin checking out its OSEMs, etc, so we could start moving it to it's new place, and figure out why it's been wonky for a while.  I ran the particle counter, and we have a factor of ~5 more particles than normal.  Kiwamu and I agreed not to open ETMY.  Since we had briefly opened the IOO and Output Optics chambers to check the X green's position on the PSL table, we immediately shut those doors.  They were probably open for ~15 minutes or so.  (Yes Steve, we should have checked before opening any doors, but at least we remembered to check at all, and the doors were only open for a few minutes rather than for a few hours.)

I attach a 24hrs trend of the particle counts, for reference.  It looks like it's been a little high for a while, but today it's really dirty in the air.

Attachment 1: ParticleCount_High_7Aug2011.png
ParticleCount_High_7Aug2011.png
  5133   Sun Aug 7 14:01:58 2011 kiwamuUpdateGreen LockingX green beam re-aligned

[Jenne / Kiwamu]

 The X green beam has been realigned to compensate the effect of the ETMX repositioning.

After the alignment we became able to lock the 00 mode with the X green beam.

 

For the alignment:

  spot position on the ETMX mirror = within ~ 1 cm. This number is strictly constrained by a homemade aluminum iris that Jamie put last Friday.

  spot position on the ITMX mirror = unknown, but looks pretty good on the CCD camera.

  spot position on the PSL table = ~ 1 mm downward from what it used to be. The horizontal alignment is perfect.

Conclusions :

  The X green beam again became a reference of the beam axis.

  The ETMX suspension tower is in a good place.

Quote from #5127

Kiwamu will work on the green alignment over the weekend.  Assuming everything works out, we'll try the same procedure on ETMY on Monday.

 

  5132   Sun Aug 7 11:10:54 2011 SureshSummaryGeneralOSEM positioning

Quote:

This OSEM placement is just the OPPOSITE of what the proper placement is.

Usually, we want to put them in so that the LED beam is vertical. This makes the OSEM immune to the optic's vertical mode.

The orientation with the horizontal LED beam makes the immunity to the side mode better, but may spoil the vertical.

In reality, neither of these assumptions is quite right. The LED beam doesn't come out straight. That's why Osamu and I found that we have to put in some custom orientations.

Also, the magnet gluings relative to the OSEM bracket centers are not perfectly aligned. So...I am saying that the OSEMs have to be oriented empirically to reduce the couplings which we want to reduce.

 

 Kiwamu suggested that since the side resonance is at a lower frequency than the bounce (~17Hz)  we ought to worry about the side more than the bounce.  If this is okay we can reposition the OSEMs to minimise this coupling. 

More over, in the current position, the OSEM s will not sense the side motion!!  So we definitely need to reposition them.  Sorry! I was being a spatz. 

  5131   Sat Aug 6 13:38:02 2011 ranaSummaryGeneralSummary of today's in-vacuum work

This OSEM placement is just the OPPOSITE of what the proper placement is.

Usually, we want to put them in so that the LED beam is vertical. This makes the OSEM immune to the optic's vertical mode.

The orientation with the horizontal LED beam makes the immunity to the side mode better, but may spoil the vertical.

In reality, neither of these assumptions is quite right. The LED beam doesn't come out straight. That's why Osamu and I found that we have to put in some custom orientations.

Also, the magnet gluings relative to the OSEM bracket centers are not perfectly aligned. So...I am saying that the OSEMs have to be oriented empirically to reduce the couplings which we want to reduce.

 

  5130   Sat Aug 6 03:10:05 2011 SureshSummaryGeneralSummary of today's in-vacuum work

The table below gives the OSEM positions as seen on the slow chanels C1:SUS-ETMX_{UL,UR,LL,LR,SD}PDMon

ETMX_OSEMs.png

Note that the side OSEM has the fast channel (OUTPUT) available and we used that to locate it.

When we began work the OSEMs were photographed so that we have a record of their locations till now.  It was difficult to get accurate estimate of the magnet offset inside the OSEM we could not see the screen on the camera while clicking.  We then took some pictures after finishing the work. These are given below

 

          Before_osem_adj.JPG         After_osem_adjustment.JPG

 

The picture of the left is from before OSEMs were moved. It can be seen that OSEMs are rotated to make sure that the magnets avoid touching the teflon sheets which hold the shadow sensors.  The picture on the right shows the positions of the OSEMs after we adjusted their positions.  This time we kept the teflon sheets vertical as shown to minimise the coupling between the Side and Axial directions.

We needed to reposition them once again after we moved the tower to the center of the table.

Pictures with more detail will be posted to the wiki later.

 

 

 

 

  5129   Fri Aug 5 22:54:29 2011 kiwamuMetaphysicsTreasureCrane crew here Monday morning

Also, according to Steve, there will be some crane guys for fixing the Y end crane issue (#5124) Monday morning.

Quote from #5128

 a film crew will be here Monday morning. They are slated to be here from 8am to 12:30pm or so. 

  5128   Fri Aug 5 20:44:26 2011 jamieMetaphysicsTreasureFilm crew here Monday morning

Just a reminder that a film crew will be here Monday morning, filming Christian Ott for some Discovery channel show.

They are slated to be here from 8am to 12:30pm or so.  They will take a couple of shots inside the lab, and the rest of the filming should be of Christian in the control room (which they will "clean up" and fit with "sexy lighting").  I will try to be here the whole time to oversee everything.

  5127   Fri Aug 5 20:37:34 2011 jamieSummaryGeneralSummary of today's in-vacuum work

[Jamie, Suresh, Jenne, Koji, Kiwamu]

After this morning's hiccup with the east end crane, we decided to go ahead with work on ETMX.

Took pictures of the OSEM assemblies, we laid down rails to mark expected new position of the suspension base.

Removed two steering mirrors and a windmill that were on the table but where not being used at all.

Clamped the test mass and moved the suspension to the edge of the table so that we could more easily work on repositioning the OSEMs.  Then leveled the table and released the TM.

Rotated each OSEM so that the parallel LED/PD holder plates were oriented in the vertical direction.  We did this in the hopes that this orientation would minimize SD -> POS coupling.

For each OSEM, we moved it through it's full range, as read out by the C1:SUS-ETMX_{UL,UR,LL,LR,SD}PDMon channels, and attempted to adjust the positions so that the read out was in the center of the range (the measured ranges, mid values, and ultimate positions will be noted in a follow-up post).  Once we were satisfied that all the OSEMs were in good positions, we photographed them all (pictures also to follow).

Re-clamped the TM and moved it into it's final position, using the rails as reference and a ruler to measure as precisely as possible :

ETMX position change: -0.2056 m = -20.56 cm = -8.09 in (away from vertex)

Rebalanced the table.

Repositioned the mirror for the ETMX face camera.

Released TM clamps.

Rechecked OSEM centering.

Unblocked the green beam, only to find that it was displaced horizontally on the test mass about half an inch to the west (-y).  Koji determined that this was because the green beam is incident on the TM at an angle due to the TM wedge.  This presents a problem, since the green beam can no longer be used as a reference for the arm cavity.  After some discussion we decided to go with the TM position as is, and to realign the green beam to the new position and relock the green beam to the new cavity.  We should be able to use the spot position of the green beam exiting the vacuum at the PSL table as the new reference.  If the green X beam exiting at the PSL table is severely displaced, we may decide to go back in and move ETMX to tweak the cavity alignment.

At this point we decided that we were done for the day.  Before closing up, we put a piece of foil with a hole in it in front of the the TM face, to use as an alignment aperture when Kiwamu does the green alignment.

Kiwamu will work on the green alignment over the weekend.  Assuming everything works out, we'll try the same procedure on ETMY on Monday.

  5126   Fri Aug 5 18:29:35 2011 JennyUpdatePSLNetwork analyzer and PD set up to measure amplitude response of PZT

Quote:

Today I placed the PDA255 photodiode on the PSL table to catch the small amount of beam power reflected by the second polarizing beam splitter in my setup. I plugged the PD output to the oscilloscope to measure the voltage output and positioned the PD such that the voltage output was maximized. At best I was able to achieve a 300 mV DC output voltage from the PD, (which seems a bit low, as the PD is specified to go from 0 to 5 V and the specifications say that the response becomes nonlinear after 10 mW/cm^2 and my beam has an intensity of approximately 5 mw/cm^2. I would therefore expect to get more beam power but after over an hour of maneuvering, 300 mV was the highest voltage output I could get).

I am planning, tomorrow afternoon, to take a measurement of the amplitude response of the PZT driving the NPRO laser. I moved the 4395 spectrum/network analyzer to near the PSL table and connected the RF output to an RF splitter. I fed one output of that into the PZT and the other output into the R port on the network analyzer. I fed the PD output into the A port. I plan to measure A/R as a function of driving frequency, sweeping from 10 Hz to 30 mHz.

I also worked to improve the mode matching of the NPRO beam coming from the AP table to the reference cavity. I drove the temperature of the NPRO at 0.100 Hz with an amplitude of 0.300 V, which Koji told me corresponds to a 1GHz change in the laser frequency. The transmission from the cavity is being monitored by a camera connected to a TV monitor, and also by a PD connected to an oscilloscope. I then repositioned the second lens in my mode matching setup in an attempt to increase the transmission peaks from the zeroth order spacial mode and decrease the transmission peaks from higher order modes. I may have improved the mode matching slightly but I was unable to improve it significantly.

The ABSL beam had been blocked so that it wouldn't enter the interferometer. I moved the block so that the beam I've been using is unblocked by the beam going to the interferometer is still blocked.

I positioned a fast lens (f=28.7mm) a little over an inch in front of the PDA255 in order to decrease the spot size incident on the PD. I adjusted the rotation angle of the half wave plate to maximize the transmitted power through the PBS to the cavity and minimize the power reflected to my PD. I then adjusted the lens potion to fix the beam on the PD. The voltage output of the PD is now 150mW, but I have the ability to increase the incident power by rotating the wave plate slightly.

Now all I need is to set up the network analyzer again to record the amplitude response to modulating the PZT from 10 Hz to 30 MHz, reduce the input voltage into the analyzer using a DC block.

  5125   Fri Aug 5 15:11:24 2011 kiwamuUpdateSUSRe: cross coupling

There is a page on the 40m wiki explaining the procedure.

  http://blue.ligo-wa.caltech.edu:8000/40m/OSEM_adjustment_proceedure

 

Additionally there are several old elogs about the cross-coupling minimization, which can be useful for us:

[1] "SUS ranking from measured data", iLog by Osamy (Aug.22 2005)

[2] "TF from position to sensors for ETMX ", iLog by Osamu (Aug.25 2005)

[3] "Further OSEM tweaking", iLog by Rana (Oct.3 2006)

Quote from #5123

We need a plan how to minimize cross coupling in the OSEMs now

 

  5124   Fri Aug 5 14:03:12 2011 steveUpdateGeneralmalfunctioning crane at ETMY

The horizontal trolley drive stopped working  at the east end this morning. It is working intermittently. In the worst case we can take the door off with the manual -Genie- lift.

I'm working with Konecrane to solve  the wormgear drive problem.

  5123   Fri Aug 5 13:51:51 2011 steveUpdateSUScross coupling

We need a plan how to minimize cross coupling in the OSEMs now

  5122   Fri Aug 5 08:08:42 2011 kiwamuSummaryGeneralplan today

Today's main mission is : adjustment of the arm length

 

   + Open the ETMX(Y) door, starting from 9:00 AM

   + Secure the ETMX(Y) test mass by tightening the earthquake stops.

   + Move the ETMX(Y) suspension closer to the door side

   + Inspect the OSEMs and take pictures before and after touching the OSEMs.

   + Level the table

   + Adjust the OSEM positions

   + Move the ETMX(Y) suspension to have designed X(Y)arm length

   + Level the table again

   + Align the ETMX(Y) such that the green beam resonate

  5121   Fri Aug 5 04:03:16 2011 KojiUpdateLSCREFL165 PD

REFL165 PD was made and tested. The characterization results are in the PDF file.

Resonance at 166.12MHz
Q of 7.3, transimpedance 667Ohm (Series Resistance = Z/Q2 = 2.5Ohm)
shotnoise intercept current = 4.3mA (i.e. current noise of 36pA/rtHz) 


As the circuit pattern had ~10nH level strain inductance, some technique was needed.

  • The diode was pushed in so as to reduce the lengths of the legs as short as possible.
  • The inductor for the resonant circuit has been located as close to the photodiode as possible
  • The other side of the inductor was needed to be bypassed by a large (0.1uF) capacitor, as the original circuit pattern (D1-L5-C33//R22) was too skinny and long.
  • C32 is also moved next to the diode.
  • The path of the photo current circuit was made thicker by Cu tapes.

Now the size of the loop for the resonant circuit is comparable with the size of SOIC-8 opamp.
(Left-Top corner of the photo)

This improved the resonant gain by factor of 8.5dB at the test with TEST INPUT. (Analyzer photo)

There is no tunable component.
The resonant freq was adjusted by a parallel inductance (270nH) to the main inductor (15nH).

P8041394.JPG

P8041395.JPG

Attachment 3: REFL165_test_110804_KA.pdf
REFL165_test_110804_KA.pdf REFL165_test_110804_KA.pdf REFL165_test_110804_KA.pdf REFL165_test_110804_KA.pdf
  5120   Fri Aug 5 02:00:09 2011 kiwamuSummaryGeneralsummary of today's invac work

[Jamie, Jenne, Suresh, Steve, Koji, Kiwamu]

We got two green beams coming out from the chambers !

 

Summary of today's invac work :

    - removed the access connector and the BS north door

    - realigned the X and Y arm to the green beams.

    - installed a HWP on the ETMY table to rotate the polarization of the green beam to P.

    - repositioned the first periscope on the BS table.

    - repositioned the second periscope on the IOO table.

    - steered some green mirrors on the IOO and OMC chamber to let the Y green beam come out to the PSL table.

    - installed a PBS in front of the first periscope to spatially overwrap two green beams.

    - adjusted the incident angle of the PBS to maximize the power of the Y green beam, which is transmitted through it.

   - steer two mirrors on the BS table to align the X green beam

   - installed two beam dumps, one is near the PBS to eliminate a ghost in the X green beam, and the other is on the back side of IPPOS/ANG pick off window.

   - closed the doors.

 

Some Notes:

 When steering the final green mirror on the OMC table, accidentally we changed the alignment of the MC incident mirror.

So the alignment of the incident beam going into MC has changed, and we haven't re-aligned it yet.

 During we were installing the PBS on the BS table, we found that the allowable incident angle for the Y beam is about ~ 55 deg, which maximizes the amount of the transmitted Y green.

Since the PBS had been considered to be 45 deg incident in our optical layout, this required several modifications in the green mirrors.

To have a clear X green beam path going into the PBS, we had to slide the PBS and periscope to the West.

The periscope is now sitting on the very edge of the BS table, and in fact ~ 20% of the bottom plate of the periscope is already sticking out.

Also since 30% of the area of the PBS's post is on a hole, which is somehow for the stack, we had to use three dog clamps instead of a folk clamps to make the contact tight.

  5119   Thu Aug 4 20:05:23 2011 NicoleUpdateSUSNew Horizontal Sliding Base Mount for TT suspension testing

In order to more-securely mount the TT supsion to the horizontal sliding base, I have made a sub-mounting plate (upon Koji's suggestion) to go in between the horizontal sliding base and the TT suspension base. I made many mistakes in this once-pristine aluminum board. I learned that using a ruler is not good enough for determining where to make holes. Upon Koji's suggestion, I have completed the mounting plate by first making a full-scale diagram on Solid Works, printing it out, and then using the diagram to determine where to make my punch holes. Thank you also to Manuel for helping me drill and to Suresh for teaching me how to use the taps!

topplateview.jpgplatform.jpg

I have been able to successfully mount the plate to the horizontal sliding platform. The TT suspension base is mounted to the front of the mounting plate (there are counter-sink screws at the front connecting the platform to the slider so that the screw heads do not obstruct the TT base). I have been able to successfully mount the TT suspension base to the mounting plate. I have also reattached the TT suspension frame to its original base (the one that I modified so that the TT could be mounted to a 1 inch pitch surface). Currently, the TT suspension is mounted to the optical table I have been working on (next to the MC-2 chamber). I am working on balancing the mirror. I am going to balance the mirror using a 670nm LED laser.

Below is a picture of the laser and the laser block I am using. After I took this photo, I have mounted the laser and the block to the optical table next to the MC-2 chamber.

 

lasersetup.jpg

I have already leveled the laser and I will plan to work on balancing the mirror tomorrow morning (my hands were shaking a lot this afternoon/evening, so I think it would be best to wait until the morning when I will be more careful). I am now going to work on the second half of my photosensor circuit box and second sensor head.

 

Please do not touch the 670nm laser on the optical table next to the MC-2 chamber! It has been leveled. Please also be careful around the optical table, since the TT suspension is mounted to the table!

  5118   Thu Aug 4 11:18:12 2011 ranaUpdateSUSsus at atm

Remember that the Oplevs are not good references because of the temperature sensitivity. The week long trend shows lots of 24 hour fluctuations.

  5117   Thu Aug 4 09:42:19 2011 KojiUpdateABSLABSL Laser shutter closed

The shutter of the ABSL laser is closed for the vent work.

  5116   Thu Aug 4 08:25:51 2011 steveUpdateSUSsus at atm

ITMX and PRM moved alot.  BS and   ITMY just a little based on oplev reference.

Attachment 1: SUSatm.png
SUSatm.png
Attachment 2: oplevsaftervent.png
oplevsaftervent.png
  5115   Thu Aug 4 01:49:08 2011 SureshUpdateIOOMC is locked

I measured the power transmitted from the PSL to the MC. It is 19mW.

The MC is now locked.  The MC Autolocker script cannot be used now since the tigger conditions are not met.  It has been disabled on the C1IOO-LOCK_MC screen.  The boost switch also is set to zero.  Increasing the boost results in MC unlocking. 

The C1:IOO-MC_RFPC_DCMON was going from 1.4 (MC Unlocked) to 0.66 (MC_locked).  I thought we ought to have a factor of ten drop in this since under high power conditions we used to have a drop of about 5.6 to 0.6.   So I adjusted the zig-zag at the end of the PSL table to improve the alignment.  It now goes from 1.4 to 0.13 when the MC is locked.   The lock is also much more stable now.  It still does not tolerate any boost though.

I checked to make sure that the beam centering on MC_REFL PD is optimal since I touched the zig-zag.  The RFPD output is now 0.7V (MC unlocked).  This matches well with the fact that we used to have 3.5V on it with the MC unlocked.  And we have cut the down the power incident on this by a factor of 5.  Because 1W -> 20mW at the PSL table  and 10% BS -> 100% Y1-...

 

 

ELOG V3.1.3-